OpenStudy (anonymous):

An insulating circular disk of radius a has a static charge. The disk rotates about its center with an angular velocity w. Find the magnetic field at the center of the disk

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

ya its zero

OpenStudy (anonymous):

actually the B field equals \[(mu _{0 w _{q}}) \div2pir\] but what I wanted to know is how to derive this result.

OpenStudy (vincent-lyon.fr):

1. divide the disc in circular strips of radius r and width dr. 2. every strip carries a charge dq = σ 2πr.dr 3. moving charge on the strip is a current di = dq/T = ω dq/2π that should help you find the magnetic field eventually. The value is indeed \(\large B_{centre}=\Large \frac{\mu_0\omega \:Q}{2\pi a} \)